OpenStudy (anonymous):

Which equation is an identity? (1 point) 7 – (9x + 3) = –9x – 4 6m – 5 = 7m + 5 – m 10p + 6 – p = 12p – 3(p – 2) 3y + 2 = 3y – 2

OpenStudy (welshfella):

simplify each side of the equation if left is same as right its an identity

OpenStudy (welshfella):

I'll do the first one for you 7 – (9x + 3) = –9x – 4 7 - 9x - 3 = -9x - 4 -9x + 4 = -9x - 4 - not an identity

OpenStudy (welshfella):

remember when you multiply by negatives: - * + = - and - * - = +